OpenStudy (anonymous):

(x^2-y^2 )dx+(y^2-2xy)dy=0

OpenStudy (anonymous):

what is the general solution to this

OpenStudy (anonymous):

Have you learned about exact equations? I wouldn't want to delve into something you're not at least somewhat familiar with.

OpenStudy (anonymous):

Uh, okay. I'll just go ahead with the exact equations method. \[x^2-y^2+\left(y^2-2xy\right)\frac{dy}{dx}=0\] For the equation to be exact, you must have \[\frac{\partial}{\partial x}\left[y^2-2xy\right]=\frac{\partial}{\partial y}\left[x^2-y^2\right]\] Indeed, you have \[-2y=-2y\] and thus the equation is exact.

OpenStudy (anonymous):

The general solution will be of the form \(\Psi(x,y)=C\), since by the chain rule we have \[\begin{align*}\frac{d}{dy}\Psi(x,y)&=\frac{d}{dy}C\\ \Psi_x+\Psi_y\frac{dy}{dx}&=0 \end{align*}\] where \(\Psi_x\) and \(\Psi_y\) denote the corresponding partial derivatives. You have \(\Psi_x=x^2-y^2\), which tells you that \[\begin{align*}\int\Psi_x~dx&=\int\left(x^2-y^2\right)~dx\\ \Psi&=\frac{1}{3}x^3-xy^2+h(y) \end{align*}\] Differentiating with respect to \(y\), you have \[\begin{align*}\Psi_y&=-2xy+\frac{dh}{dy}\\ y^2-2xy&=-2xy+\frac{dh}{dy}\\ y^2~dy&=dh\\ \int y^2~dy&=\int dh\\ \frac{1}{3}y^3+C&=h(y)\end{align*}\] Thus, the solutions is \[\Psi=\frac{1}{3}x^3-xy^2+\frac{1}{3}y^3+C~~\Rightarrow~~\frac{1}{3}x^3-xy^2+\frac{1}{3}y^3=C^*\]

OpenStudy (anonymous):

Is the equation second order and linear?

OpenStudy (anonymous):

Second order? No, the highest-order derivative is one. Linear? Also no, there's a factor of \(y^2\) that makes it non-linear.

OpenStudy (anonymous):

do you know how to do this? Verify that y(x)=-1/14+12e^14x is a solution of the ordinary differential equation y^'=14y+1. Note you do not need to solve the ordinary differential equation.

OpenStudy (anonymous):

Plug in the given \(y\) into the equation. If equality holds, then it's a solution.

OpenStudy (anonymous):

so I plug the top into the bottom one?

OpenStudy (anonymous):

Yes. \[y=-\frac{1}{14}+12e^{14x}~~\Rightarrow~~y'=168e^{14x}\] Plug these in

OpenStudy (anonymous):

That's all I have to do? so it is verified?